OpenStudy (anonymous):

what is the best way to prove 7^k + 2^(2k+1) is always divisible by 3

OpenStudy (anonymous):

\[7^{k} + 2^{2k+1}\] I've been trying by induction, but can't quite get it.Any help much appreciated

OpenStudy (anonymous):

first i show that for n =1 its true i.e 7 + 2^3 = 15, divisible by 3 then assume true for k, it should be true for \[7^{k+1} + 2^{2(k+1)+1)}\] however what to do next?

OpenStudy (anonymous):

first, the base case: k = 1: 7^1 + 2^(2+1) = 7 + 8 = 15, which is divisible by 3 now assume 7^k + 2^(2k+1) is divisible by 3 ie for some integer b, 7^k + 2^(2k + 1) = 3b now we need to show that for some arbitrary integer d: 7^(k+1) + 2^(2(k+1) + 1) = 3d 7^(k+1) + 2^(2(k+1) + 1) = 7*7^k + 2^(2k+1+2) = 7*7^k + 4*2^(2k+1) = 3*7^k + 4*7^k + 4*2^(2k+1) = 3*7^k + 4[7^k + 2^(2k+1)] = 3*7^k + 4*3b = 3*7^k + 12b = 3[7^k + 4b] Since 7^k is an integer, and 4b is an integer, then 7^k + 4b is integer hence we have proven that 7^(k+1) + 2^(2(k+1) + 1) = 3d, for some arbitrary integer d completes our proof of induction

OpenStudy (nikvist):

\[7\equiv 1\enspace\,(mod\enspace3)\quad\Rightarrow\quad 7^k\equiv 1^k\enspace\,(mod\enspace3)\quad\Rightarrow\quad 7^k\equiv 1\enspace\,(mod\enspace3)\] \[2^2\equiv 1\enspace\,(mod\enspace3)\quad\Rightarrow\quad 2^{2k}\equiv 1^k\enspace\,(mod\enspace3)\quad\Rightarrow\quad 2^{2k}\equiv 1\enspace\,(mod\enspace3)\] \[\Rightarrow\quad 2^{2k+1}\equiv 2\enspace\,(mod\enspace3)\] Finally, \[7^k+2^{2k+1}\equiv 1+2\equiv 0\enspace(mod\enspace 3)\] proof without induction

OpenStudy (anonymous):

fantastic - thanks for the help

OpenStudy (anonymous):

yeah, nikvist's proof is very elegant if you want to use modular arithmetic